A French investor buys 230 shares of Teck for $14,950 ($65 per share). Over the course of a year, Teck goes up by $8.65.

a. If there is a 10 percent gain in the value of the dollar versus the euro, what will be the total percentage return to the French investor? (Do not round intermediate calculations. Round the final answer to 2 decimal places.)

Total percentage return            %

b. Now assume the stock increases by $10 but the dollar decreases by 10 percent versus the euro. What will be the total percentage return to the French investor? (Round the final answer to 2 decimal places.)

Total percentage return            %
Answer #1

a.

Return on Stock = 8.65/65 = 13.31%

Return to French Investor = (1.1331)(1.10) - 1

Return to French Investor = 24.64%

b.

Return on Stock = 10/65 = 15.38%

Return to French Investor = (1.1538)(1 - 0.10) - 1

Return to French Investor = 26.92%